OpenStudy (anonymous):

Systems of equations with the same slope and different y intercepts have no solution. Sometimes Always Never

OpenStudy (anonymous):

6.03] The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 840! If three times the number of one-step equations minus twice the number of two-step equations is equal to 870, how many one-step equations auditioned today? 330 510 520 660 Also this...

OpenStudy (pounce0129):

sometimes

OpenStudy (anonymous):

Can you explain why

OpenStudy (anonymous):

Please?

OpenStudy (pounce0129):

if it has the same slope then the lines are parallel and this means they never cross. when you solve a system you find the point at which the lines intersect and since they never intersect there isn't going to be a solution all the time

OpenStudy (anonymous):

Oh..okay thank you! I need to pass this course this year! :) Do you know the answer to... 6.03] The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 840! If three times the number of one-step equations minus twice the number of two-step equations is equal to 870, how many one-step equations auditioned today? 330 510 520 660

OpenStudy (anonymous):

I'll be your fan if you want...:)

OpenStudy (pounce0129):

you could try to look it up on google, i just started algebra this year

OpenStudy (anonymous):

Do you understand that question? Oh...you started algebra! Me too! Online.

OpenStudy (cwrw238):

create system of equations in x and y where x = one-step and y = two-step equations.

OpenStudy (pounce0129):

all i have got so far is 3x-y=870 where x is one-step equations and y is two-step equations

OpenStudy (cwrw238):

x + y = 840

OpenStudy (anonymous):

Yeah...but this qestion is weird! :( Can somone tell me the answer and tell me how to work it out?

OpenStudy (cwrw238):

3x - 2y = 870

OpenStudy (pounce0129):

x - 1-step y - 2-step x + y = 840 3x - 2y = 870 --Use substitution-- 3(840 - y) - 2y = 870 --Solve for y-- y = 330 --plug y in the first equation-- x + 330 = 840 --solve for x-- x = 510 found it here: http://answers.yahoo.com/question/index?qid=20100529231454AAAiKbc

OpenStudy (anonymous):

So the answer is 510?

OpenStudy (pounce0129):

yeppers

OpenStudy (anonymous):

Ok...thank you! :) I got it now! :)

OpenStudy (cwrw238):

x + y = 840 3x-2y = 870 multiply first equation by 3: 3x + 3y = 2520 3x - 2y = 870 subtract 5y = 1650 y = 330 = 2 step one step x = 840-330 = 510

OpenStudy (anonymous):

Ok, I become both of you guy's fans! :)

OpenStudy (cwrw238):

oh ok i used the elimination method pounce used the substitution method

OpenStudy (anonymous):

6.03] What is the value of the x variable in the solution to the following system of equations? 3x + 2y = 8 2x + 3y = 2 4 −2 x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system Do you guys know this problem?

OpenStudy (pounce0129):

i used the google method actually

OpenStudy (anonymous):

True True! :)

OpenStudy (cwrw238):

pounce - are u sure about the first question? - sometimes?

OpenStudy (anonymous):

What I hate is all these essay questions you have to type up! :(

OpenStudy (pounce0129):

i put sometimes?? i meant always!!

OpenStudy (anonymous):

Ok, I got that one! :) Thanks! You guys are very helpful!:)

OpenStudy (cwrw238):

ok

OpenStudy (pounce0129):

for the first one its always not sometimes my mistake

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

Okay, I fixed it just now! 6.03] Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 2x + 5y = −3 −15x − 5y = −10

OpenStudy (pounce0129):

http://answers.yahoo.com/question/index?qid=20110728095221AAj2cv9 i feel so helpful!!

OpenStudy (anonymous):

Yes, with google (narrows eyes) JK! You are very helpful!

OpenStudy (cwrw238):

elimination is best method for this you only need to add the 2 equations and y is eliminated

OpenStudy (cwrw238):

-13x = -13 x = 1

OpenStudy (anonymous):

Thank you! You guys are very helpful! Anything you need help with?

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

:)

OpenStudy (pounce0129):

i need help woth one thing: is f(x)=-2/1 the same thing as y=-2/1?

OpenStudy (anonymous):

Is it mulitple choice or an essay question?

OpenStudy (pounce0129):

neither its asking to write an equation in function notation form and i got -2/1

OpenStudy (anonymous):

Oh, can you put the original ? here I can see it. :)

OpenStudy (anonymous):

It kinda is the same because they both equal -2/1, but also different sice you have to find 2 equations for one of them and 1 for the other.

OpenStudy (pounce0129):

thanks